# Autonomous Navigation Of Quadrupeds Using Coverage Path Planning

### v1

@@ III-B Path Planning
Algorithm 3 A method to find the nearest leaf vertex $\boldsymbol{v}_{leaf,target}$ from a selected of leaf nodes $\mathcal{V}^{\prime}_{leaf}$ to a given source $\boldsymbol{x}_{0}$ given the Euclidean distances.
This is iterated until all leaf vertices have been added and $\mathcal{P}$ is complete. Nevertheless, because the original distance between every two waypoints is approximately equal to the map’s resolution, this leads to significant time spent on scanning with minimal displacement. Knowing that the robot has a large scanning range capability allowing to scan distances greater than the map resolution, we can increase the distance between every two waypoints to some arbitrary distance $D$ by reducing $P$ by a factor of $D/R$ such that $D\geq R$. This is done by selecting every $D/R$ element in $\mathcal{P}$.
@@ III-C State Machine
To achieve autonomy that enables the robot to visit the waypoints and perform a scanning procedure in succession, an activity diagram is formulated as seen in Figure 4(a). We can group two or more activities as a singular state if the transitions are consecutive without interruptions e.g. a decision node. This state will perform all activities in their respective order. On the other hand, each decision node becomes a state that checks if its control variable reached a specified threshold. Each state has a conditionless trigger that automatically transitions the current state to the next specified state at the end of its action.
With a finite state machine (FSM), we can achieve the desired autonomous navigation based on the triggers by handling the sequence of the actions. The implemented state machine diagram can be seen in Figure 4(b). Each state is explained as follows:
  • In State Load Map, an existing 2D navigation map is loaded onto SLAM and allows the robot to be localized with respect to the map’s metadata (meter-per-pixel resolution, width and height both in pixels, and origin as $(x,y)$ in meters), and the robot’s concurrent surroundings through 2D laser scans. After this map is loaded, it is then analyzed to generate a list of waypoints, each in $(x,y,\psi)$ coordinate, for the robot to visit using map reader. An optimal route is then planned to visit all of these waypoints which reorders the original list using path planner. Finally, it automatically transitions to the next state, Check Waypoints.
@@
  • In State Check Destination, the state machine determines if the robot is already at the current destination with some acceptable offset. This is done by determining if the Euclidean distance between the robot’s 2D pose $\mathbf{x}$ and a destination’s pose $\mathbf{x}_{\text{des}}$, is less than arbitrary tolerance $\boldsymbol{\delta}$. Note that $\mathbf{x},\mathbf{x}_{\text{des}},\boldsymbol{\delta}\in\mathbb{R}^{3}$ since the poses are defined in the $x$- and $y$-axes, and the yaw orientation $\psi$ with respect to /map frame. If this comparison is true, it transitions to State Scan. Otherwise, it transitions to State Move.
  • In State Move, it actuates the robot to move towards the desired destination. Once the navigation is successful or timed out, it transitions back to State Check Destination. However, if it detects an interruption by a human operator via the joystick controller, it cancels the ongoing action and instantaneously transitions to State Manual Control.
  • In State Scan, the robot performs the procedure to scan the local environment. The scanning procedure is detailed in Appendix -C. Once it is completed, it transitions to State Check Waypoints.
@@ IV Experiments and Results
The Go2 Edu robot was initially manually controlled via a wireless controller to generate a 2D navigation map using the SLAM module. The experiment was then conducted and the whole process was repeated over five trials. The SLAM module’s map resolution was set to $0.10$ meters. The point cloud buffer time was set to 0.25 seconds. The laser scan minimal and maximal range was set to 0.50 meters and 30.0 meters, respectively. The smoothing standard deviation was set to 3. The crispification threshold was set to 128. The erosion kernel $K$ was set to 10 x 10 pixels. The waypoint-to-waypoint distance $D$ in path planning was set at 1.00 m. The robot’s maximum x-, y- and, yaw velocity were set to 1.00 m/s, 0.50 m/s, and 0.80 rad/s. The 2D navigation position and orientation tolerance were set to 0.05 m and 0.08 rad, respectively, with a timeout of 10.0 s. The scanning procedure module was disabled to demonstrate the navigational tasks.
@@ IV-A SLAM, Map Reader and Path Planning
To evaluate the time efficiency of the map reader, we record the duration required to process each map across five trials, considering a total of $N=100$ instances. We vary the weights in terms of the dimension per map. To facilitate a clearer representation in the plot, we use the product of the dimensions, $H\times W$ in pixels. As can be seen in Figure 7(a), we can see that the average time taken for map reader ranges from 2.34 ms to 2.70 ms over the five trials with an average standard deviation of about 0.30 ms. We can also see that the mean time increases by 22.0 ns for every additional pixel in the map. For instance, with a map with size $H\times W=10^{6}$ pixels, it has an expected mean time $T_{read}=23.6$ ms. To remain within a maximum time taken of 1 s, the map should have not be larger than the size of $H\times W=45.4\times 10^{6}$ pixels. Assuming a square map with map resolution of 0.10 m, this maximum map has a length of 673.4 m. Hence, our proposed map reader is fast for relatively small to large map sizes (ranging up to $45.4\times 10^{6}$ pixels) within a lead time of $T_{read}=1.0$ s, especially when the map reader is only run once to produce the waypoints given the map’s geometry.
@@ IV-B Navigational Performance
3 82.61 287.2 5.40 Significant map drift occurred after 3 minutes such that the remaining 8 waypoints were unreachable (displaced into the occupied space).
4 100.0 322.3 5.40 Robot stalled at Waypoints 6, 10-11, 18, 23, 35, 37-39 due to replanning.
5 48.72 160.1 5.40 Significant map drift occurred after 2 minutes such that the remaining 20 waypoints were unreachable (displaced into the occupied space).
To determine the robot’s navigational performance, we evaluate it according to the two aforementioned metrics. This can be seen in Table II. In addition, in Figure  9, the time taken for each waypoint over all trials can be seen. We can see that the robot is able to reach $86.5$% of the total waypoints over all trials with a median time taken of $5.38$ s per waypoint. As seen in Figure 9, in an average time of $8.525$ s with a standard deviation of $11.4$ s. We can see several outliers where the robot took ranging from 10.0 seconds to a few minutes to complete the navigation to the adjacent waypoint, for instance at Waypoint 17 in Trial 1. This can be explained due to Nav2 trying to create a local path that neatly ends within the desired set tolerances in position and orientation, and the fact that the robot requires a minimal input velocity in order to move, it overshoots beyond the tolerance and Nav2 has to replan the local path. This can result in the robot being stuck indefinitely, and requires assistance from a human operator. This means that for some of these outliers ranging from 30 s, human assistance is eventually brought into action. Such that the robot is correctly placed in the desired position and orientation within the set tolerance.
Figure 9: The time taken in seconds for the robot to reach each possible unique waypoint over five trials. Gray dots describe waypoints that the robot has not been able to reach. The outliers occurred due to Nav2 getting stuck at replanning such that the navigation terminates within narrow desired tolerances.
@@ V Conclusion
Coverage path planning allows mobile robots such as quadrupeds to explore the whole environment. This is especially useful for applications such as surveillance, inspection, and search and rescue. Nevertheless, limited works are done on autonomous navigation using coverage path planning on quadrupeds. Therefore, we developed an open-source framework using ROS 2 that enables an Unitree Robotics Go2 quadruped to navigate such that it visits every corner using the prior 2D navigation map, autonomously. It utilizes a map reader to extract a graph of 2D waypoints using the topological skeleton of the map, and a path planner to create an efficient path with respect to time and the starting position. A state machine is used to iterate over the ordered list of waypoints and navigate them in succession.
@@
Future work will focus on expanding the proposed method with real-time autonomous coverage exploration in unknown environments in 3D, verify the algorithm on Unitree Robotics Go2-W for its enhanced mobility and adaptability in challenging terrains, examine and improve the map drift in SLAM, and compare to existing 2D coverage path planning methods.

### v2

@@ III-B Path Planning
Algorithm 3 A method to find the nearest leaf vertex $\boldsymbol{v}_{\text{leaf,target}}$ from a set of selected leaf nodes $\mathcal{V}^{\prime}_{\text{leaf}}$ to a given source $\boldsymbol{x}_{0}$ given the Euclidean distances.
This is iterated until all leaf vertices have been added and $\mathcal{P}$ is complete. Nevertheless, because the original distance between every two waypoints is approximately equal to the map’s resolution, this leads to significant time spent on scanning with minimal displacement. Considering that the robot has a large scanning range capability of 30 m, which allows it to scan distances greater than the map resolution, we can increase the distance between every two waypoints to some arbitrary distance $D$ by reducing $P$ by a factor of $D/R$ where $D\geq R$. This is done by selecting every $D/R$-th element in $\mathcal{P}$.
@@ III-C State Machine
To achieve autonomy that enables the robot to visit the waypoints and perform a scanning procedure in succession, an activity diagram is formulated as seen in Figure 4(a). We can group two or more activities as a singular state if the transitions are consecutive without interruptions, e.g., a decision node. This state will perform all activities in their respective order. On the other hand, each decision node becomes a state that checks if its control variable has reached a specified threshold. Each state has a conditionless trigger that automatically transitions the current state to the next specified state at the end of its action.
With a finite state machine (FSM), we can achieve the desired autonomous navigation based on the triggers by handling the sequence of actions. The implemented state machine diagram can be seen in Figure 4(b). Each state is explained as follows:
  • In State Load Map, an existing 2D navigation map is loaded onto SLAM and allows the robot to be localized with respect to the map’s metadata (meter-per-pixel resolution, width and height both in pixels, and origin as $(x,y)$ in meters), and the robot’s concurrent surroundings through 2D laser scans. After this map is loaded, it is then analyzed to generate a list of waypoints, each in $(x,y)$ coordinates, for the robot to visit using the map reader. An optimal route is then planned to visit all of these waypoints, which reorders the original list using the path planner. Finally, it automatically transitions to the next state, Check Waypoints.
@@
  • In State Check Destination, the state machine determines if the robot is already at the current destination with some acceptable offset. This is done by determining if the Euclidean distance between the robot’s 2D pose $\boldsymbol{x}$ and a destination’s pose $\boldsymbol{x}_{\text{des}}$, is less than or equal to a set tolerance $\boldsymbol{\delta}$, where $\boldsymbol{x},\boldsymbol{x}_{\text{des}},\boldsymbol{\delta}\in\mathbb{R}^{3}$ since the poses are defined in the $x$- and $y$-axes, and the yaw orientation $\psi$ with respect to /map frame. If the condition is true, it transitions to State Scan. Otherwise, it transitions to State Move. It is worth noting that the path planner does not account for the orientation $\psi$. However, this orientation is likely a necessary requirement of the scanning procedure, ensuring that the robot is aligned with the desired orientation, as described in Supplementary Material C.
  • In State Move, it actuates the robot to navigate towards the desired destination. Navigation is considered successful when the Euclidean distance between the robot’s current pose $\boldsymbol{x}$ and the target destination $\boldsymbol{x}_{\text{des}}$ is less than or equal to a predefined threshold $\delta$. Alternatively, if the navigation process exceeds a specified timeout duration $T_{\text{timeout}}$, it is also terminated. In either case, the system transitions back to the Check Destination state. However, if the system detects an interruption by a human operator via the joystick controller mid-operation, the system promptly cancels the ongoing action and immediately transitions to State Manual Control.
  • In State Scan, the robot performs the procedure to scan the local environment, which is detailed in Supplementary Material C. Once it is completed, it transitions to State Check Waypoints.
@@ IV Experiments and Results
The Go2 Edu robot was initially manually controlled via a wireless controller to generate a 2D navigation map using the SLAM module. The experiment was then conducted and the whole process was repeated over five trials. The SLAM module’s map resolution was set to $0.10$ meters. The point cloud buffer time was set to 0.25 seconds. The laser scan’s minimal and maximal range was set to 0.50 meters and 30.0 meters, respectively. The smoothing standard deviation was set to 3. The crispification threshold was set to 128. The erosion kernel $K$ was set to 10 x 10 pixels. The waypoint-to-waypoint distance $D$ in path planning was set at 1.00 m. The robot’s maximum x-, y- and, yaw velocities were set to 1.00 m/s, 0.50 m/s, and 0.80 rad/s, respectively. The 2D navigation position and orientation tolerance were set to 0.05 m and 0.08 rad, respectively, with a navigation timeout of 10.0 s to replan. The scanning procedure module was disabled to demonstrate the navigational tasks.
@@ IV-A SLAM, Map Reader and Path Planning
To evaluate the time efficiency of the map reader, we record the duration required to process each map across five trials, considering a total of $N=100$ instances. We vary the weights in terms of the dimension per map. To facilitate a clearer representation in the plot, we use the product of the dimensions, $H\times W$ in pixels. As can be seen in Figure 7(a), we can see that the average time taken for the map reader ranges from 2.34 ms to 2.70 ms over the five trials with an average standard deviation of about 0.30 ms. We can also see that the mean time increases by 22.0 ns for every additional pixel in the map. For instance, with a map with size $H\times W=10^{6}$ pixels, it has an expected mean time $T_{\text{read}}\approx 23.6$ ms.
To remain within a maximum time taken of 1.0 s, the map should not be larger than the size of $H\times W=45.4\times 10^{6}$ pixels. Assuming a square map with a map resolution of 0.10 m, this maximum map has a length of 673.4 m. Hence, our proposed map reader is fast for relatively small to large map sizes (ranging up to $45.4\times 10^{6}$ pixels) within a lead time of $T_{\text{read}}=1.0$ s, especially when the map reader is only run once to produce the waypoints given the map’s geometry.
@@
To stay within a lead time of 1.0 s, the number of waypoints should not be larger than $1.219\times 10^{5}$. The bottleneck appears if the map resolution $R$ becomes small, as it increases the number of pixels per meter in a map. This can substantially increase the number of waypoints by a factor of $R/R^{\prime}$, where $R^{\prime}$ is the new map resolution and $R^{\prime}<R$. Nonetheless, the path planner can be considered fast for a small to large number of waypoints, given a reasonable map resolution $R$. It is also worth noting that it also runs only once to produce the necessary path.
@@ IV-B Navigational Performance
3 82.61 287.2 5.40 Significant map drift occurred after 3 minutes, such that the remaining 8 waypoints were unreachable (displaced into the occupied space).
4 100.0 322.3 5.40 Robot stalled at Waypoints 6, 10, 18, 23, 28, 35, 37-39 due to replanning.
5 48.72 160.1 5.40 Significant map drift occurred after 2 minutes, such that the remaining 19 waypoints were unreachable as they displaced into the occupied space.
To determine the robot’s navigational performance, we evaluate it according to the two aforementioned metrics. This can be seen in Table II. In addition, in Figure  9, the time taken for each waypoint over all trials can be seen. We can see that the robot is able to reach $86.5$% of the total waypoints over all trials with a median time taken of $5.38$ s per waypoint. As seen in Figure 9, this was done in the average time of $8.525$ s with a standard deviation of $11.4$ s.
We can see several outliers where the robot took ranging from 10.0 seconds to a few minutes to complete the navigation to the adjacent waypoint, for instance, at Waypoint 17 in Trial 1. This can be explained due to Nav2 trying to create a local path that neatly ends within the desired set tolerances in position and orientation, and the fact that the robot requires a minimal input velocity in order to move, it overshoots beyond the tolerance and Nav2 has to replan the local path. This can result in the robot being stuck indefinitely, and requires assistance from a human operator. This means that for some of these outliers, ranging from 30 s, human assistance is eventually brought into action. Such that the robot is correctly placed in the desired position and orientation within the set tolerance.
Figure 9: The time taken in seconds for the robot to reach each possible unique waypoint over five trials. Gray dots describe waypoints that the robot has not been able to reach. The outliers occurred due to Nav2 getting stuck at replanning, such that the navigation terminates within narrow desired tolerances, or due to map drift.
@@ V Conclusion
Coverage path planning allows mobile robots such as quadrupeds to explore the whole environment. This is especially useful for applications such as surveillance, inspection, and search and rescue. Nevertheless, limited work has been done on autonomous navigation using coverage path planning on quadrupeds. Therefore, we developed an open-source framework using ROS 2 that enables a Unitree Robotics Go2 quadruped to navigate such that it visits every corner using the prior 2D navigation map, autonomously. It utilizes a map reader to extract a graph of 2D waypoints using the topological skeleton of the map, and a path planner to create an efficient path with respect to time and the starting position. A state machine is used to iterate over the ordered list of waypoints and navigate them in succession.
@@
Compared to the state-of-the-art methods that use time-constrained planning [23, 28], which do not enable variable time of task execution, our approach is not constrained based on a predefined time. While a time-constrained approach is useful when mission time is predefined, the variable-time approach provides a more adaptable solution when operating with limited knowledge. Unlike alternative state-of-the-art approaches that enable time-variable exploration [24], our approach is more computationally efficient due to the use of a morphological technique to extract the topological skeleton of the map. Even though the experimental conditions were different, a rough comparison of the map generation magnitude based on the results from each paper shows an order-of-magnitude difference (milliseconds vs microseconds).
Nevertheless, the proposed method is primarily suited for known and static environments, rendering it unsuitable for real-world applications with irregular and moving obstacles within the environment, and varying elevations. Future work will therefore focus on extending the proposed method to incorporate real-time autonomous coverage exploration in unknown 3D environments, particularly under diverse environmental conditions and in the presence of irregular and dynamic obstacles. Additionally, the influence of variations in key parameters will be systematically investigated. Moreover, the study will also investigate and seek to improve drift issues commonly encountered in SLAM. Lastly, the proposed approach will be systematically compared to existing 2D coverage path planning methods to evaluate its relative performance and advantages.